ubmit a paper which is 2-3 pages in length (no more than 4-pages), exclusive of the reference page. The paper should be double spaced in Times New Roman (or its equivalent) font which is no greater than 12 points in size. The paper should cite at least three sources in APA format. One source can be your textbook. In this paper, please discuss the following case study. In doing so, explain your approach to the problem, support your approach with references, and execute your approach. Provide an answer to the case study’s question with a recommendation. You are the owner of a parasailing company that is expanding operations to a new beachfront location, and you need to prepare a 3-year analysis for the bank that may loan you the funds to purchase your boat and parasailing equipment. A lot of business is done on a referral basis, where a company pays a fee to a 3rd party to send them customers. However, because of your well-established reputation, you already have received requests for “flights” to be scheduled as soon as you open the new location. Therefore, you expect to break-even the first year but must calculate the number of flights needed. You also need to determine the new break-even point in Year 2 if the location allows referrals, which you believe will cost on average about 2% of the sales price overall. Finally, you need to determine the volume needed to have $10,000 in profit in Year 3. The following information is available: Sales price per flight $175 Estimated loan payment per month $350 Fuel costs per flight $100 Full-time scheduler salary $2,500 per month Boat crew per flight $30 $500 per month dock fee and use of a small office on a pier Requirements: Calculate the Year 1 break-even quantity, contribution margin, and contribution margin ratio. Explain how the values were determined. Calculate the Year 2 break-even quantity, break-even sales, and contribution margin ratio. Explain how the values were determined. Determine the number of flights (units) needed to retain a profit of $10,000 in Year 3, assuming the company does allow for referrals. Recommend if the bank should issue the loan. Superior papers will: Perform all calculations correctly. Articulate the approach to solving the problem. Explain the relationship of the costs to the concept of contribution margin. Discuss any limitations of the data, including what may be missing. Conclude on whether the bank should issue the loan.
Answer #1
YEAR 1 BREAKEVEN QUANTITY
A Sales Price per flight $175
Variable Costs:
B Fuel Cost $100
C Boat Crew $30
D=B+C Total Variable Costs $130
E=A-D Unit Contribution margin $45
Fixed Costs:
F Loan Repayment $350
G Salary of scheduler $2,500
H Dock fee and use of office $500
I=F+G+H Total Monthly Fixed Costs $3,350
J=I*12 Annual Fixed Cost $40,200
Break Even Quantity=Fixed Cost/Unit Contribution
K=J/E YEAR 1 BREAKEVEN QUANTITY 894 (40200/45) (Rounded to higher whole number)
Contribution Margin $40,230 (894*45)
Contribution Margin Ratio=Contribution Margin/Sales
Contribution Margin Ratio 0.25714286 (45/175)
Contribution Margin Ratio 25.71%
YEAR 2
Sales Price per Flight= $175
Cost of referral=2%*175= $3.50
Net Sales Revenue=175-3.5 $171.50
Variable Cost per unit $130
Unit Contribution Margin $41.50
Contribution Margin Ratio 0.24198251 (41.50/171.50)
Contribution Margin Ratio 24.20%
YEAR 2 BREAKEVEN QUANTITY 969 (40200/41.50)
Contribution Margin $40,214 (969*41.50)
Break even Sales=969*175 $169,575
YEAR 3 REQUIRED PROFIT $10000 (WITHOUT REFERRALS)
Break even number of flights without referrals 894
Contribution margin per flight $45.00
Number of flights above Break even number                222 (10000/45)
Total Number of flights required =894+222             1,116